### v2

# GAICo: A Deployed and Extensible Framework for Evaluating Diverse and Multimodal Generative AI Outputs

Abstract
The rapid proliferation of Generative AI (GenAI) into diverse, high-stakes domains necessitates robust and reproducible evaluation methods. However, practitioners often resort to ad-hoc, non-standardized scripts, as common metrics are often unsuitable for specialized, structured outputs (e.g., automated plans, time-series) or holistic comparison across modalities (e.g., text, audio, and image). This fragmentation hinders comparability and slows AI system development. To address this challenge, we present GAICo (Generative AI Comparator): a deployed, open-source Python library that streamlines and standardizes GenAI output comparison. GAICo provides a unified, extensible framework supporting a comprehensive suite of reference-based metrics for unstructured text, specialized structured data formats, and multimedia (images, audio). Its architecture features a high-level API for rapid, end-to-end analysis, from multi-model comparison to visualization and reporting, alongside direct metric access for granular control. We demonstrate GAICo’s utility through a detailed case study evaluating and debugging complex, multi-modal AI Travel Assistant pipelines. GAICo empowers AI researchers and developers to efficiently assess system performance, make evaluation reproducible, improve development velocity, and ultimately build more trustworthy AI systems, aligning with the goal of moving faster and safer in AI deployment. Since its release on PyPI in Jun 2025, the tool has been downloaded over 13K times, across versions, by Aug 2025, demonstrating growing community interest.
Deployed Librarypypi.org/project/GAICo
Codegithub.com/ai4society/GenAIResultsComparator
Documentationai4society.github.io/projects/GenAIResultsComparator
Introduction
The adoption of Generative AI (GenAI), particularly Large Language Models (LLMs), has been transformative across industries. However, this rapid integration has outpaced the development of standardized, accessible, and domain-appropriate evaluation tools. While standard Natural Language Processing (NLP) metrics like BLEU (Papineni et al. 2002) and ROUGE (Lin 2004) are useful, they are often insufficient for structured outputs like AI plans or time-series forecasts. In the absence of a unified tool, developers create bespoke evaluation scripts, leading to a fragmented and non-reproducible landscape that hinders the comparison and iterative improvement of AI systems.
Common scenarios for usage of GAICo are evaluating an LLM’s output along a range of metrics (Figure 1) and output of multiple LLMs across a single metric (notebooks ‘example-1’ and ‘example-2’ in Table 3). Consider the practical challenge of building a significantly more complex composite “AI Travel Agent” that uses an orchestrator LLM to generate an itinerary and specialist models to create corresponding images and audio summaries (System Pipelines in Figure 2). Evaluating different combinations of these components requires writing separate, ad-hoc scripts for each data modality. This process is slow, error-prone, and makes it difficult to pinpoint the root cause of a poor output: was it a flawed prompt from the orchestrator or poor execution by a specialist model? This exemplifies the need for a unified tool that can efficiently handle complex, multi-modal comparisons.
Figure 1: The multi-modal GAICo workflow. The framework processes answers from multi-modal (text, image, audio) AI models, computes pairwise similarity scores ($s_{kl}$), and constructs several outputs: raw data reports, visualizations, and pass/fail assessments against a threshold $\delta$ (any distance function, or conversely, 1 - similarity metric). See metrics in Table 5.
Figure 2: Illustration of GAICo for a composite AI system case study. (Right) A user uses LLMs to get results tailored to their needs and then analyzes them with GAICo. They start the process in three parallel pipelines that generate multi-modal outputs (JSON plan, image, audio). The user then performs a two-part evaluation: (a) Plan Coherence, comparing the JSON outputs of all pipelines against a single Baseline Plan Reference (derived from Pipeline A’s output) and (b) Modality Quality, comparing each pipeline’s generated image and audio against per-pipeline references. These per-pipeline references are generated by feeding the same prompts/scripts from the respective pipeline’s JSON output into baseline specialist generators (from Pipeline A). (Left) The output radar plot shows image and audio fidelity relative to references derived from each pipeline’s prompts. For more examples, see Table 3.
In response, we introduce GAICo (Generative AI Comparator), a deployed and extensible open-source Python library built to address these challenges. GAICo provides a unifying framework that makes the application of diverse, reference-based metrics simple and reproducible across different data types (modalities). By providing a consistent interface, GAICo directly supports the iterative workflows common in applied AI development. According to a PyPI package statistics source (PyPI Stats 2025), the library has been downloaded over 13,000 times in just over 2 months (Jun-Aug 2025), demonstrating its widespread and growing community interest.
Our contributions, aligned with the “Tools and Methodologies for Moving Faster and Safer” track, are: (1) A Unified and Extensible Framework: A general framework for comparing GenAI outputs, with an object-oriented design that allows seamless integration of metrics for text, structured data (AI plans, time-series), and multimedia. (2) A Streamlined Practitioner Workflow: The Experiment class, which simplifies the end-to-end evaluation of complex AI systems by abstracting score calculation, visualization, and reporting into a few lines of code. (3) Demonstrated Utility for Debugging Composite AI: A case study on AI Travel Assistants (Figure 2) showcasing GAICo’s effectiveness in performing nuanced, multi-stage evaluations to isolate performance issues. (4) A Deployed and Maintained Tool: A public tool on PyPI with comprehensive documentation, optional dependencies to manage installation size, and a robust testing suite to lower the barrier to adoption.
Related Works
General-Purpose AI/NLP Evaluation Libraries
General-purpose AI and Natural Language Processing (NLP) evaluation libraries provide foundational metrics for various tasks. For instance, the Hugging Face evaluate library (Wolf et al. 2019) offers a wide array of metrics for NLP and related tasks, providing a standardized compute method for individual metrics and integrating well within the Hugging Face ecosystem. Similarly, Scikit-learn (Pedregosa et al. 2011) provides fundamental metrics for classification, regression, and clustering, adaptable for certain aspects of text evaluation (e.g., cosine similarity on TF-IDF vectors). Libraries like NLTK (Bird, Klein, and Loper 2009) and SpaCy (Honnibal et al. 2020) offer foundational text processing capabilities and some basic metrics (e.g., BLEU in NLTK).
While indispensable, these libraries primarily offer individual text metrics and lack a unified framework for evaluating diverse, multi-modal GenAI outputs or streamlining multi-model comparisons. This fragmentation forces developers to manually integrate disparate tools. GAICo directly addresses this issue by providing a single, extensible API that unifies evaluation across diverse data types, significantly reducing development overhead and improving reproducibility.
A key differentiator for GAICo is its design choice to decouple evaluation from LLM inference, positioning it as a post-hoc, reproducible comparison framework. This contrasts with general-purpose libraries like Hugging Face evaluate by offering a more integrated workflow and specialized metrics, and with end-to-end frameworks, e.g., Ragas (Es et al. 2024), DeepEval (Ip and Vongthongsri 2025), that are tightly coupled with LLM APIs for “LLM-as-a-judge” evaluations. While these integrated frameworks are powerful, their coupling can introduce complexities like API costs, rate-limits, and nondeterminism. GAICo’s focus on pre-generated outputs simplifies the evaluation pipeline, making it faster and more reliable for developers focused on comparative analysis. The Experiment class, discussed in Section 3The GAICo Framework, further streamlines this by offering a high-level abstraction for multi-model comparison, plotting, and reporting. A detailed comparison of these approaches is provided in Section 1B of the Appendix.
Domain-Specific Evaluation Approaches
Evaluation in specialized AI domains often relies on custom tools, leading to fragmented practices. GAICo bridges this gap by integrating relevant metrics into a single, consistent framework.
Automated Planning: Evaluation in automated planning typically focuses on properties like plan length, cost, or state-space coverage, often requiring specialized planners or validators (e.g., PDDL parsers). However, when LLMs generate planning sequences, the output might not strictly adhere to formal planning languages, necessitating metrics that compare the content and order of actions rather than strict plan validity. GAICo addresses this with PlanningLCS and PlanningJaccard metrics, which are designed to robustly compare action sequences, including concurrent actions, directly from LLM outputs. These are inspired by foundational work on measuring plan diversity and similarity (Srivastava et al. 2006, 2007) and sequence analysis (Hirschberg 1977), adapted specifically for GenAI evaluation contexts.
Time-Series Analysis: Standard time-series evaluation metrics include Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), or R-squared for forecasting accuracy; however, when comparing generated time-series shapes or sequences of values from LLMs, metrics like Dynamic Time Warping (DTW) (Berndt and Clifford 1994) become crucial for handling variations in speed or phase. GAICo integrates TimeSeriesDTW and TimeSeriesElementDiff to provide robust, normalized comparisons of time-series data, regardless of their textual representation.
Multimedia (Image/Audio): Evaluation of generated images and audio often involves specialized libraries and metrics. For images, common metrics include Structural Similarity Index (SSIM) (Wang et al. 2004), Peak Signal-to-Noise Ratio (PSNR) (Sara, Akter, and Uddin 2019), or perceptual hashes (Farid 2021). For audio, metrics like Signal-to-Noise Ratio (SNR) (Yuan et al. 2019), Perceptual Evaluation of Speech Quality (PESQ) (Rix et al. 2001), or Short-Time Objective Intelligibility (STOI) (Taal et al. 2010) are used. While underlying libraries (e.g., scikit-image (Van der Walt et al. 2014), Pillow (Clark 2015), librosa (McFee et al. 2015), scipy (Virtanen et al. 2020)) provide these functionalities, GAICo unifies their application within its framework. The rise of multi-modal generative models, such as Meta’s ImageBind (Girdhar et al. 2023), further underscores the need for unified evaluation tools capable of assessing outputs across diverse modalities. This allows for consistent, normalized comparison of multimedia outputs alongside text and structured data, bridging the gap between general LLM evaluation and highly specialized domain-specific analysis.
GAICo’s contribution lies in integrating these diverse, domain-specific metrics into a single, extensible library, accelerating development and improving the reliability of composite AI systems.
The GAICo Framework
The GAICo framework is architected to provide a robust, extensible, and user-friendly solution for evaluating GenAI outputs. Its design balances a high-level, streamlined workflow with the flexibility required for granular analysis, centered on three components: an extensible BaseMetric class, a comprehensive metric library, and a high-level Experiment class. A UML class diagram for GAICo  can be viewed in Figure 7 (Section 3 in the Appendix).
Core Concepts: BaseMetric and Extensibility
At the heart of GAICo’s architecture is the BaseMetric abstract class, which establishes a consistent foundation for all metrics. It mandates that every metric, regardless of data modality, implements a single core method: calculate(generated_texts, reference_texts). This simple interface is key to the framework’s flexibility, transparently handling various input formats (single items, lists, NumPy arrays) for efficient batch processing. This architecture makes GAICo highly extensible: a developer can introduce a new metric by simply inheriting from BaseMetric and implementing the calculate() method, immediately integrating it into the GAICo ecosystem.
Comprehensive Metric Library
GAICo provides a carefully designed library of metrics spanning textual analysis, structured data, and multimedia (audio and image). This comprehensive approach addresses the reality that modern GenAI applications produce complex, multi-modal outputs that traditional text-only tools cannot adequately assess. The library includes n-gram-based (BLEU, ROUGE), text similarity (Jaccard, Cosine), and semantic (BERTScore) metrics. For structured data, it offers specialized metrics for automated planning sequences (PlanningLCS) and time-series (TimeSeriesDTW). For multimedia, it integrates established metrics for image (SSIM, PSNR) and audio (SNR, Spectrogram Distance) quality. Detailed descriptions of each metric are available in Section 3D of the Appendix.
The Streamlined Experiment Class
The Experiment class is a high-level API designed to streamline the entire evaluation workflow. It orchestrates the common task of comparing multiple model outputs against a single reference, encapsulating score calculation, visualization, and reporting into a few lines of code. Its compare() method automates the generation of scores, publication-ready plots (bar or radar), application of quality thresholds, and export of comprehensive CSV reports. By abstracting away boilerplate code, the Experiment class directly improves developer productivity and enables rapid, reproducible benchmarking.
Deployment and Maintainability
GAICo is a professionally engineered tool deployed on the Python Package Index (PyPI) for easy installation via pip. To balance functionality with accessibility, it uses optional dependencies (e.g., pip install ‘gaico[bertscore]’), minimizing the installation footprint. The project’s quality is ensured through a comprehensive test suite (pytest (Krekel et al. 2004)), continuous integration (CI) for automated testing and deployment, and pre-commit hooks (Sottile 2014) for code consistency. Extensive documentation (MkDocs (Christie 2014)), including runnable Jupyter notebooks and a full API reference, lowers the barrier to adoption for the applied AI community.
Case Study: Evaluating Composite
AI Travel Assistants
To demonstrate GAICo’s utility in a realistic scenario, we evaluated three distinct, composite AI systems tasked with creating a multi-modal travel itinerary. This study highlights how GAICo enables a nuanced evaluation crucial for debugging and improving such systems. An overview of the case study is demonstrated in Figure 2. The code for the case study can be found in notebook number 7, ‘case_study.ipynb’, in Table 3.
Figure 3: Radar plots generated by GAICo comparing pipeline performance across various metrics. (Right) Modality Generation Quality, assessing the specialist models’ fidelity against references generated from their own pipeline’s prompts. Each axis is a metric, and each line represents a pipeline’s averaged score. (Left) Plan Coherence, showing the relative strengths of orchestrator LLMs against a universal human-curated reference.
Setup and Methodology
Our goal was to assess AI systems built from different combinations of state-of-the-art generative models. Each system, or “pipeline”, consists of an orchestrator LLM that generates a travel plan in JSON format, which in turn provides prompts for downstream, specialized models for image and audio generation.
Master Orchestrator Prompt
All orchestrator LLMs received the same detailed prompt, designed to elicit a 3-day Paris itinerary in a structured JSON format, including fields for planning sequences, time-series, image prompts, and audio prompts (latter two for the specialist models). The full prompt is included in Section 2 of the Appendix.
We constructed three pipelines using different combinations of orchestrator LLMs and specialist image/audio models (Pipeline A: OpenAI-centric, Pipeline B: Open-Source, Pipeline C: Google-centric). Each pipeline received an identical master prompt to generate a 3-day Paris itinerary in a structured JSON format, which included prompts for the downstream models.
A key challenge in composite systems is attributing failure. To address this, we designed a two-part evaluation strategy, seamlessly executed using GAICo: (1) Plan Coherence: We evaluated the core travel planning ability of the orchestrator LLMs by comparing the generated JSON content (text, plans, budgets) from all pipelines against a single high-quality reference plan (from Pipeline A). (2) Modality Generation Quality: We evaluated the fidelity of the specialist models by comparing each pipeline’s generated image and audio against a reference created using that pipeline’s own generated prompts. This isolates the specialist model’s execution quality from the orchestrator’s prompt-generation quality.
Table 1: Plan Coherence Results. Scores are averaged over three days, comparing each pipeline’s orchestrator output against a single baseline reference. Higher scores between Pipeline B and C are highlighted.
Pipeline ROUGE-L BERTScore-F1 PlanningLCS PlanningJaccard TimeSeriesDTW
Pipeline A 1.000 1.000 1.000 1.000 1.000
Pipeline B 0.190 0.599 0.095 0.083 0.122
Pipeline C 0.222 0.613 0.137 0.117 0.367
Table 2: Modality Generation Quality Results. Scores are averaged over three days, comparing each pipeline’s specialist model outputs to a high-quality reference generated from the same prompt. Higher scores between Pipeline B and C are highlighted.
Pipeline Image SSIM Image AvgHash Image HistMatch Audio SNR Audio SpecDist
Pipeline A 1.000 1.000 1.000 1.000 1.000
Pipeline B 0.276 0.646 0.642 0.249 0.261
Pipeline C 0.347 0.766 0.528 0.247 0.260
Results and Analysis
The quantitative results, summarized in Tables 1 and 2 and visualized in Figure 3, demonstrate GAICo’s value in accelerating AI development and improving system safety.
Orchestrator Performance (Plan Coherence)
As shown in Table 1, Pipeline A’s orchestrator (GPT-5), serving as the high-quality baseline, expectedly achieved perfect scores (1.000) across all plan coherence metrics. In contrast, both Pipeline B (Llama 4) and Pipeline C (Gemini 2.5 Pro) showed significantly lower performance, indicating substantial deviations from the baseline plan. As shown in the radar plot in Figure 3 (left), Pipeline C generally outperformed Pipeline B across these metrics, suggesting a better ability to approximate the reference budget forecast.
Specialist Model Performance (Modality Generation Quality)
The modality generation quality results (Table 2) reveal the execution fidelity of the specialist models against their respective high-quality references. Pipeline A’s specialist models achieved perfect scores (1.000) as their outputs were used to define the modality-specific references. For image generation, Pipeline C generally outperformed Pipeline B across Image SSIM and Image AverageHash, indicating stronger structural and perceptual similarity to the baseline. However, Pipeline B showed a slightly higher Image HistogramMatch score, suggesting its color distribution was closer to the baseline. This highlights the value of multi-metric evaluation for a nuanced understanding.
Focusing on the audio metrics, both Pipeline B and Pipeline C produced audio summaries with notably lower Audio SNR and Audio SpectogramDist compared to Pipeline A. This indicates that while their orchestrators might provide valid scripts, their specialist audio models are not yet capable of matching the fidelity of the baseline OpenAI TTS. Figure 3 (right) visually represents these performance differences.
Analysis
This case study vividly demonstrates how GAICo facilitates “moving faster and safer” in the development of complex AI systems.
Faster Development Cycles Without GAICo, this multi-modal evaluation would require numerous, disparate scripts for JSON parsing, text analysis, plan validation, and multimedia comparison, followed by manual aggregation. GAICo streamlined this entire process into a single, unified workflow. This enables rapid iteration, allowing developers to swap components (e.g., an orchestrator LLM) and immediately receive quantitative feedback on the impact across all modalities, drastically reducing the time and maintenance costs associated with benchmarking.
Improved Reliability and Safer Deployment The two-part evaluation, enabled by GAICo, is critical for enhancing system reliability. By separating “Plan Coherence” from “Modality Generation Quality”, we can precisely diagnose failures. For instance, the results show that Pipeline B’s poor performance stems from both a weak orchestrator (low scores in Table 1) and less-than-optimal specialist models (Table 2). This granular insight allows for targeted optimization, improving the orchestrator’s planning logic or replacing a specific specialist model, which is essential for building trustworthy systems and preventing incidents caused by underperforming components. This directly contributes to the goal of producing better AI solutions by enabling a more rigorous and efficient development process.
Demonstrated Use Cases and Documentation
To maximize ease of use and adoption, key factors in a tool’s real-world impact, GAICo is supported by extensive documentation and a comprehensive suite of example notebooks. These resources demonstrate the library’s versatility across a range of applications, from simple quick-start guides to advanced, domain-specific evaluations in finance, planning, and multimedia. Table 3 summarizes the available examples, which are all publicly accessible in our code repository.
Table 3: Summary of GAICo Example Notebooks. Those referenced in paper are emphasized. All available at:
https://github.com/ai4society/GenAIResultsComparator/tree/main/examples.
Notebook Name Description
Quickstart Examples
01. case_study.ipynb Full code for the composite AI travel assistant evaluation presented in this paper.
02. example-1.ipynb Evaluating multiple models using a single metric via the calculate() method.
03. example-2.ipynb Evaluating a single model across multiple metrics using their calculate() methods.
04. example-audio.ipynb Evaluating AI-generated audio using specialized audio metrics.
05. example-image.ipynb Evaluating AI-generated images using specialized image metrics.
06. example-structured_data.ipynb Comparing text-based metrics with specialized metrics for time-series and planning.
07. quickstart.ipynb A simple, quickstart workflow using GAICo’s Experiment module.
Intermediate Examples
08. example-DeepSeek.ipynb Evaluating DeepSeek for Point of View (POV) document analysis.
09. example-audio_data.ipynb Advanced audio analysis for TTS and music generation tasks.
10. example-election.ipynb Evaluating models on sensitive election-related questions.
11. example-finance.ipynb Evaluating models on finance domain questions by iterating over a dataset.
12. example-planning.ipynb Evaluation of various travel plans using planning-specific metrics.
13. example-recipes.ipynb Evaluating models on recipe generation with parallelization via joblib.
14. example-timeseries.ipynb In-depth evaluation and perturbation of time-series data.
Advanced Examples
15. example-llm_faq.ipynb Comparing various LLM responses (Phi, Mixtral, etc.) on a FAQ dataset.
16. example-threshold.ipynb Exploring default and custom thresholding techniques for LLM evaluation.
17. example-viz.ipynb Demonstrations of creating custom visualizations for evaluation results.
Conclusion
In this section, we highlight the unique role of GAICo in evaluating LLM outputs in the context of prior work. Then, we discuss how it fits into the role of “moving faster and safer” in AI system development.
GAICo and Its Novelty
GAICo addresses a critical gap in GenAI evaluation by providing the first unified, extensible, and deployed framework for reproducible assessment across diverse data modalities. Recent studies have found that even with access to the original source code, many ML experiments cannot be reproduced, mainly due to missing details in documentation and inconsistent evaluation procedures (Lopresti and Nagy 2021). GAICo replaces this current landscape of ad-hoc, non-comparable scripts with a standardized approach, integrating metrics for text, structured data (planning, time-series), and multimedia. This standardization is GAICo’s core novelty; it ensures that once a metric is chosen, its application and reporting are consistent, directly improving the quality and productivity of the evaluation process. As demonstrated in our case study, this unified approach enables nuanced debugging, such as distinguishing orchestrator faults from specialist model failures, which is crucial for enhancing the reliability of complex AI systems.
GAICo as a Deployed Tool
GAICo is engineered as a practical utility to help the community move faster and safer. Its design prioritizes ease of use and adoption through a simple PyPI installation, optional dependencies to manage installation size, comprehensive documentation, and a suite of runnable examples. This accessibility allows teams to accelerate development cycles; the streamlined Experiment class replaces disparate scripts with a single, consistent API, improving developer productivity. The framework enables a safer deployment of AI by facilitating precise failure analysis. By distinguishing a poor plan from a poor execution, GAICo improves system reliability, reduces long-term maintenance costs, and allows teams to mitigate weaknesses before they become incidents. Ultimately, GAICo is a deployed tool designed to have a significant impact by enabling the applied AI community to produce better AI solutions through rigorous, standardized, and efficient evaluation.
Limitations and Future Work
Despite its strengths, GAICo has some limitations in its current scope. First, the current metric suite, while broad, primarily focuses on reference-based comparisons and does not yet include metrics for fairness, bias, toxicity, or operational aspects like latency and computational cost. Furthermore, current visualizations are static plots (bar, radar). There is no built-in interactive dashboard or web-based UI for dynamic exploration of results. Finally, while supporting structured data like planning and time-series, the library does not yet offer general-purpose comparison for arbitrary complex structured data (e.g., nested JSONs, knowledge graphs) or multi-turn conversational outputs.
We now discuss how one may extend the limitations in the future. A key priority is to incorporate metrics for fairness, bias, toxicity, and operational statistics, providing developers with a more holistic toolkit for building safer and more responsible AI systems. GAICo’s extensible design makes this easy (see UML class diagram, Figure 7). Future work also includes exploring developing interactive dashboards or deeper integration with MLOps platforms (like MLflow (Zaharia et al. 2018)) to enable dynamic exploration and tracking of evaluation results. Lastly, further work would also focus on adding general-purpose comparison capabilities for complex structured data (e.g., JSON diffing, graph similarity) and extending metrics to handle multi-turn conversational outputs.
Acknowledgments
This work is partially supported by NSF Awards #2454027, NAIRR250014, and Faculty Award by JP Morgan Research.
Appendix A Appendix
This appendix provides supplementary material to support the main paper. It includes a detailed comparison of our evaluation paradigm with existing tools, all the necessary artifacts for reproducing the AI Travel Assistant case study, and comprehensive descriptions of the diverse metrics implemented in our framework.
  • Section 1B: A comparative analysis of GAICo against other evaluation paradigms, highlighting its unique features.
  • Section 2C: All artifacts for the AI Travel Assistant case study, including the orchestrator prompt and supplementary visualizations.
  • Section 3D: Detailed descriptions of the textual, structured, and multimedia evaluation metrics available in GAICo.
Appendix B Comparison with Other Evaluation Paradigms
To better situate GAICo within the broader landscape of AI evaluation tools, Table 4 provides a detailed comparison against two dominant paradigms: general-purpose metric libraries, exemplified by Hugging Face evaluate, and end-to-end, LLM-integrated frameworks like Ragas and DeepEval. The table highlights GAICo’s unique focus on post-hoc, reproducible evaluation of diverse data types (including specialized structured and perceptual multimedia metrics) and its streamlined workflow for multi-model comparative analysis, distinguishing it from libraries focused on standard NLP/CV tasks and frameworks tightly coupled with LLM inference.
Table 4: Comparison of GAICo with other AI Evaluation Paradigms
Feature / Approach GAICo Hugging Face evaluate LLM-Integrated Frameworks
Evaluation Paradigm Post-hoc, reference-based comparison Post-hoc computation of individual metrics End-to-end, often reference-free (LLM-as-a-judge)
of pre-generated outputs
Primary Scope Text, Structured Data (Planning/TS), General NLP & CV task metrics RAG-specific metrics (e.g., Faithfulness, Relevancy)
Perceptual Multimedia (e.g., BLEU, Accuracy, COCO)
High-level Workflow ✓ (Experiment class for multi-model ✗ (Focus on single metric compute calls) ✓ (Orchestrates generation & evaluation pipeline)
analysis, plotting, reporting)
Decoupled from Generation ✓ (Ensures reproducibility, avoids ✓ (Operates on pre-generated data) ✗ (Tightly coupled with LLM inference APIs)
API costs/latency during eval)
Built-in Visualization ✓ (Plots for comparative analysis) ✗ (Requires external plotting libraries) ✗ Varies (Some offer dashboards, not general plotting)
Extensibility ✓ (Inherit BaseMetric) ✓ (Inherit Metric) ✓ (Often supports custom metrics/prompts)
Appendix C Case Study Artifacts
This section contains the material necessary for the GAICo AI Travel Assistant case study (Section 4Case Study: Evaluating Composite AI Travel Assistants). The remaining scripts can be found in the GAICo repository.
Orchestrator Prompt
The orchestrator prompt shown in Figure 4 is the single system instruction fed to the top-level LLM that steers the entire AI Travel Assistant pipeline. It fulfils three roles: (i) it elicits a structured JSON itinerary that can be parsed reliably by downstream code; (ii) it embeds sub-prompts (image_prompt and audio_script) that seed the specialist image and TTS generators, thereby keeping all modalities semantically aligned; and (iii) it enforces output-format contracts that make automated evaluation with GAICo possible.
Orchestrator Prompt You are a sophisticated AI Travel Assistant Orchestrator. Your task is to generate a travel plan based on the user’s request. Your output MUST be a single, valid JSON object and nothing else. Do not include any introductory text, explanations, or markdown formatting like ```json. The user’s request is: “A 3-day Paris trip for a first-time visitor on a moderate budget.” Please generate a JSON object with a root key “trip_plan” which is an array of objects, one for each day. Each daily object must contain the following keys: “day”: (Integer) The day number. “day_plan_text”: (String) A descriptive, paragraph-long summary of the day’s activities. “day_plan_sequence”: (String) A comma-separated sequence of actions representing the plan, e.g., ”visit(location), eat(restaurant), see(landmark)”. “day_budget_euros”: (Number) An estimated budget in Euros for the day. “image_prompt”: (String) A detailed, artistic, and descriptive prompt for an image generation model, capturing the essence of the day’s main landmark. “audio_script”: (String) A concise, engaging 15-second summary of the day’s activities, written as a script for a text-to-speech model. Now, generate the complete JSON object for the 3-day trip.
Figure 4: Prompt used by the orchestrator LLM. The prompt enforces JSON output, supplies slot-level requirements for each day, and instructs downstream specialist generators (image and TTS) via embedded sub-prompts.
Supplementary Bar Charts
To complement the radar plots shown in the main paper, Figures 5 and 6 provide a metric-by-metric breakdown for the three pipelines evaluated in the AI Travel Assistant case study.
Figure 5: Plan-coherence view.
The five bars in each subplot correspond to the specialised textual or structured metrics discussed. A score of 1.0 denotes perfect agreement with the baseline reference plan produced by Pipeline A. The chart shows that Pipeline C consistently outperforms Pipeline B, hinting that the models in its pipeline improved temporal consistency.
Figure 6: Modality-quality view.
Here, we isolate the image and audio specialist generators. Because each pipeline’s own outputs serve as its reference, a perfect score of 1.0 appears only for Pipeline A, which supplies those references. Lower bars for Pipelines B and C quantify perceptual deviations. Pipeline C’s higher Image_AverageHash score versus Pipeline B indicates better content preservation despite equal SSIM degradation. Both alternatives exhibit similar audio issues, as reflected by overlapping Audio_SNR and Audio_SpectrogramDist values.
Together, the two figures pinpoint that most performance regressions stem from plan-generation weaknesses rather than the downstream modality specialists, thereby guiding future remediation work.
Appendix D Detailed Metric Library Descriptions
This section provides detailed descriptions of the metrics available in GAICo, summarized in Table 5. For an overview UML class diagram of GAICo, view Figure 7.
Textual Evaluation Framework
Our textual metrics encompass three complementary evaluation paradigms.
  • N-gram based metrics include BLEU for translation-quality assessment, ROUGE variants (rouge1, rouge2, rougeL) for summarization, and JSDivergence Jensen-Shannon Divergence for vocabulary distribution analysis.
  • Text similarity metrics provide orthogonal perspectives. JaccardSimilarity measures set-based word overlap, CosineSimilarity captures semantic relationships through TF-IDF vector space modeling, LevenshteinDistance quantifies edit-based differences, and SequenceMatcherSimilarity provides character-level analysis.
  • Semantic Similarity is implemented via BERTScore, which leverages pre-trained BERT embeddings to compute contextual similarity, addressing the limitations of surface-level metrics.
Structured Data Evaluation
We developed specialized metrics for critical structured data applications.
  • Planning sequence evaluation addresses automated planning outputs. PlanningLCS implements longest common subsequence matching that respects temporal ordering and handles concurrent actions. PlanningJaccard focuses on action set completeness, ignoring ordering.
  • Time-series evaluation tackles temporal data. TimeSeriesElementDiff implements a weighted comparison of keyed time-points. TimeSeriesDTW leverages Dynamic Time Warping to compare value sequences, effectively handling phase-shifted data.
Multimedia Evaluation Capabilities
  • Audio evaluation includes AudioSNRNormalized, which provides a normalized Signal-to-Noise Ratio on a [0, 1] scale, and AudioSpectrogramDistance, which evaluates similarity through frequency-time domain analysis using STFT.
  • Image evaluation provides a suite of perceptual and structural metrics. SSIM measures perceived quality through structural information, luminance, and contrast. PSNR quantifies pixel-wise fidelity. AverageHash generates perceptual hashes for content similarity, and HistogramMatch evaluates color distribution similarity.
Figure 5: Per-metric comparison of plan coherence across the three pipelines. Higher is better.
Figure 6: Per-metric comparison of modality-generation quality (images and audio) across the three pipelines.
Table 5: GAICo’s Comprehensive Metric Library
Category Metric Description Primary Use Cases
Textual BLEU N-gram precision with brevity penalty Translation, text generation quality
ROUGE Recall-oriented n-gram and LCS matching Summarization, content coverage
JSDivergence Symmetric vocabulary distribution comparison Bias detection, style consistency
JaccardSimilarity Set-based word overlap analysis Keyword extraction evaluation
CosineSimilarity TF-IDF vector space comparison Document semantic similarity
LevenshteinDistance Character-level edit distance Error analysis, text correction
SequenceMatcher Python difflib-based similarity Code generation, precise matching
BERTScore BERT embedding-based semantic similarity Deep semantic understanding
Structured PlanningLCS LCS with concurrent action support Action sequence evaluation
PlanningJaccard Set-based action completeness Plan coverage assessment
TimeSeriesElementDiff Weighted temporal data comparison Forecasting evaluation
TimeSeriesDTW Dynamic time warping alignment Phase-shifted sequence analysis
Audio AudioSNRNormalized Signal-to-noise ratio assessment Audio quality evaluation
AudioSpectrogramDistance STFT-based frequency analysis Audio generation evaluation
Image SSIM Structural similarity index Image quality assessment
PSNR Peak signal-to-noise ratio Image fidelity evaluation
AverageHash Perceptual hash for content similarity Duplicate detection, content matching
HistogramMatch Color distribution similarity analysis Global color profile comparison
Figure 7: UML Class Diagram showcasing GAICo’s capabilities and broad multi-media support.

### v3

@@
 GAICo empowers AI researchers and developers to efficiently assess system performance, make evaluation reproducible, improve development velocity, and ultimately build more trustworthy AI systems, aligning with the goal of moving faster and safer in AI deployment.
-Since its release on PyPI in Jun 2025, the tool has been downloaded over 13K times, across versions, by Aug 2025, demonstrating growing community interest.
+Since its release on PyPI in Jun 2025, the tool has been downloaded over 15K times, across versions, by Nov 2025, demonstrating growing community interest.
 Deployed Librarypypi.org/project/GAICo
 Codegithub.com/ai4society/GenAIResultsComparator
-Documentationai4society.github.io/projects/GenAIResultsComparator
+Documentationai4society.github.io/projects/GenAIResultsComparatsor
 Introduction
@@
 However, this rapid integration has outpaced the development of standardized, accessible, and domain-appropriate evaluation tools.
-While standard Natural Language Processing (NLP) metrics like BLEU (Papineni et al.
-2002) and ROUGE (Lin 2004) are useful, they are often insufficient for structured outputs like AI plans or time-series forecasts.
+While standard Natural Language Processing (NLP) metrics like BLEU (papineni-etal-2002-bleu) and ROUGE (lin-2004-rouge) are useful, they are often insufficient for structured outputs like AI plans or time-series forecasts.
 In the absence of a unified tool, developers create bespoke evaluation scripts, leading to a fragmented and non-reproducible landscape that hinders the comparison and iterative improvement of AI systems.
+Figure 1: The multi-modal GAICo workflow.
+The framework processes answers from multi-modal (text, image, audio) AI models, computes pairwise similarity scores ($s_{kl}$), and constructs several outputs: raw data reports, visualizations, and pass/fail assessments against a threshold $\delta$.
+(any distance function, or conversely, 1similarity metric1-$\text${similarity metric}).
+In response, we introduce GAICo (Generative AI Comparator), a deployed and extensible open-source Python library built to address these challenges.
+GAICo provides a unifying framework that makes the application of diverse, reference-based metrics simple and reproducible across different data types (modalities).
+By providing a consistent interface, GAICo directly supports the iterative workflows common in applied AI development.
+According to a PyPI package statistics source (pepy_gaico), the library has been downloaded over 13,000 times in just over 2 months (Jun-Aug 2025), demonstrating its widespread and growing community interest.
 Common scenarios for usage of GAICo are evaluating an LLM’s output along a range of metrics (Figure 1) and output of multiple LLMs across a single metric (notebooks ‘example-1’ and ‘example-2’ in Table 2).
@@
 This exemplifies the need for a unified tool that can efficiently handle complex, multi-modal comparisons.
-Figure 1: The multi-modal GAICo workflow.
-The framework processes answers from multi-modal (text, image, audio) AI models, computes pairwise similarity scores ($s_{kl}$), and constructs several outputs: raw data reports, visualizations, and pass/fail assessments against a threshold $\delta$ (any distance function, or conversely, 1 - similarity metric).
-See metrics in Table 5.
 Figure 2: Illustration of GAICo for a composite AI system case study.
@@
 For more examples, see Table 2.
-In response, we introduce GAICo (Generative AI Comparator), a deployed and extensible open-source Python library built to address these challenges.
-GAICo provides a unifying framework that makes the application of diverse, reference-based metrics simple and reproducible across different data types (modalities).
-By providing a consistent interface, GAICo directly supports the iterative workflows common in applied AI development.
-According to a PyPI package statistics source (PyPI Stats 2025), the library has been downloaded over 13,000 times in just over 2 months (Jun-Aug 2025), demonstrating its widespread and growing community interest.
 Our contributions, aligned with the “Tools and Methodologies for Moving Faster and Safer” track, are: (1) A Unified and Extensible Framework: A general framework for comparing GenAI outputs, with an object-oriented design that allows seamless integration of metrics for text, structured data (AI plans, time-series), and multimedia.
@@
 General-purpose AI and Natural Language Processing (NLP) evaluation libraries provide foundational metrics for various tasks.
-For instance, the Hugging Face evaluate library (Wolf et al.
-2019) offers a wide array of metrics for NLP and related tasks, providing a standardized compute method for individual metrics and integrating well within the Hugging Face ecosystem.
-Similarly, Scikit-learn (Pedregosa et al.
-2011) provides fundamental metrics for classification, regression, and clustering, adaptable for certain aspects of text evaluation (e.g., cosine similarity on TF-IDF vectors).
-Libraries like NLTK (Bird, Klein, and Loper 2009) and SpaCy (Honnibal et al.
-2020) offer foundational text processing capabilities and some basic metrics (e.g., BLEU in NLTK).
+For instance, the Hugging Face evaluate library (wolf2019huggingface) offers a wide array of metrics for NLP and related tasks, providing a standardized compute method for individual metrics and integrating well within the Hugging Face ecosystem.
+Similarly, Scikit-learn (scikit-learn) provides fundamental metrics for classification, regression, and clustering, adaptable for certain aspects of text evaluation (e.g., cosine similarity on TF-IDF vectors).
+Libraries like NLTK (bird2009natural) and SpaCy (Honnibal_spaCy_Industrial-strength_Natural_2020) offer foundational text processing capabilities and some basic metrics (e.g., BLEU in NLTK).
 While indispensable, these libraries primarily offer individual text metrics and lack a unified framework for evaluating diverse, multi-modal GenAI outputs or streamlining multi-model comparisons.
@@
 A key differentiator for GAICo is its design choice to decouple evaluation from LLM inference, positioning it as a post-hoc, reproducible comparison framework.
-This contrasts with general-purpose libraries like Hugging Face evaluate by offering a more integrated workflow and specialized metrics, and with end-to-end frameworks, e.g., Ragas (Es et al.
-2024), DeepEval (Ip and Vongthongsri 2025), that are tightly coupled with LLM APIs for “LLM-as-a-judge” evaluations.
+This contrasts with general-purpose libraries like Hugging Face evaluate by offering a more integrated workflow and specialized metrics, and with end-to-end frameworks, e.g., Ragas (es2024ragas), DeepEval (Ip_deepeval_2025), that are tightly coupled with LLM APIs for “LLM-as-a-judge” evaluations.
 While these integrated frameworks are powerful, their coupling can introduce complexities like API costs, rate-limits, and nondeterminism.
@@
 GAICo addresses this with PlanningLCS and PlanningJaccard metrics, which are designed to robustly compare action sequences, including concurrent actions, directly from LLM outputs.
-These are inspired by foundational work on measuring plan diversity and similarity (Srivastava et al.
-2006, 2007) and sequence analysis (Hirschberg 1977), adapted specifically for GenAI evaluation contexts.
-Time-Series Analysis: Standard time-series evaluation metrics include Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), or R-squared for forecasting accuracy; however, when comparing generated time-series shapes or sequences of values from LLMs, metrics like Dynamic Time Warping (DTW) (Berndt and Clifford 1994) become crucial for handling variations in speed or phase.
+These are inspired by foundational work on measuring plan diversity and similarity (srivastava2006finding; srivastava2007domain) and sequence analysis (hirschberg1977algorithms), adapted specifically for GenAI evaluation contexts.
+Time-Series Analysis: Standard time-series evaluation metrics include Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), or R-squared for forecasting accuracy; however, when comparing generated time-series shapes or sequences of values from LLMs, metrics like Dynamic Time Warping (DTW) (berndt1994using) become crucial for handling variations in speed or phase.
 GAICo integrates TimeSeriesDTW and TimeSeriesElementDiff to provide robust, normalized comparisons of time-series data, regardless of their textual representation.
 Multimedia (Image/Audio): Evaluation of generated images and audio often involves specialized libraries and metrics.
-For images, common metrics include Structural Similarity Index (SSIM) (Wang et al.
-2004), Peak Signal-to-Noise Ratio (PSNR) (Sara, Akter, and Uddin 2019), or perceptual hashes (Farid 2021).
-For audio, metrics like Signal-to-Noise Ratio (SNR) (Yuan et al.
-2019), Perceptual Evaluation of Speech Quality (PESQ) (Rix et al.
-2001), or Short-Time Objective Intelligibility (STOI) (Taal et al.
-2010) are used.
-While underlying libraries (e.g., scikit-image (Van der Walt et al.
-2014), Pillow (Clark 2015), librosa (McFee et al.
-2015), scipy (Virtanen et al.
-2020)) provide these functionalities, GAICo unifies their application within its framework.
-The rise of multi-modal generative models, such as Meta’s ImageBind (Girdhar et al.
-2023), further underscores the need for unified evaluation tools capable of assessing outputs across diverse modalities.
+For images, common metrics include Structural Similarity Index (SSIM) (wang2004image), Peak Signal-to-Noise Ratio (PSNR) (sara2019image), or perceptual hashes (farid2021overview).
+For audio, metrics like Signal-to-Noise Ratio (SNR) (yuan2019signal), Perceptual Evaluation of Speech Quality (PESQ) (rix2001perceptual), or Short-Time Objective Intelligibility (STOI) (taal2010short) are used.
+While underlying libraries (e.g., scikit-image (van2014scikit), Pillow (clark2015pillow), librosa (mcfee2015librosa), scipy (2020SciPy-NMeth)) provide these functionalities, GAICo unifies their application within its framework.
+The rise of multi-modal generative models, such as Meta’s ImageBind (girdhar2023imagebind), further underscores the need for unified evaluation tools capable of assessing outputs across diverse modalities.
 This allows for consistent, normalized comparison of multimedia outputs alongside text and structured data, bridging the gap between general LLM evaluation and highly specialized domain-specific analysis.
@@
 For multimedia, it integrates established metrics for image (SSIM, PSNR) and audio (SNR, Spectrogram Distance) quality.
-Detailed descriptions of each metric are available in Section 3D of the Appendix.
+Figure 3: Radar plots generated by GAICo comparing pipeline performance across various metrics.
+(Right) Modality Generation Quality, assessing the specialist models’ fidelity against references generated from their own pipeline’s prompts.
+Each axis is a metric, and each line represents a pipeline’s averaged score.
+(Left) Plan Coherence, showing the relative strengths of orchestrator LLMs against a universal human-curated reference.
 The Streamlined Experiment Class
@@
 To balance functionality with accessibility, it uses optional dependencies (e.g., pip install ‘gaico[bertscore]’), minimizing the installation footprint.
-The project’s quality is ensured through a comprehensive test suite (pytest (Krekel et al.
-2004)), continuous integration (CI) for automated testing and deployment, and pre-commit hooks (Sottile 2014) for code consistency.
-Extensive documentation (MkDocs (Christie 2014)), including runnable Jupyter notebooks and a full API reference, lowers the barrier to adoption for the applied AI community.
+The project’s quality is ensured through a comprehensive test suite (pytest (pytest)), continuous integration (CI) for automated testing and deployment, and pre-commit hooks (precommit_2025) for code consistency.
+Extensive documentation (MkDocs (mkdocs_2025)), including runnable Jupyter notebooks and a full API reference, lowers the barrier to adoption for the applied AI community.
 Case Study: Evaluating Composite
@@
 The code for the case study can be found in notebook number 7, ‘case_study.ipynb’, in Table 2.
-Figure 3: Radar plots generated by GAICo comparing pipeline performance across various metrics.
-(Right) Modality Generation Quality, assessing the specialist models’ fidelity against references generated from their own pipeline’s prompts.
-Each axis is a metric, and each line represents a pipeline’s averaged score.
-(Left) Plan Coherence, showing the relative strengths of orchestrator LLMs against a universal human-curated reference.
+Plan Coherence Metrics Modality Generation Metrics
+Pipeline $\text{ROUGE-L}_{\textit{T}}$ $\text{BERTScore-F1}_{\textit{T}}$ $\text{LCS}_{\textit{Pl}}$ $\text{Jaccard}_{\textit{Pl}}$ $\text{DTW}_{\textit{TS}}$ $\text{SSIM}_{\textit{I}}$ $\text{AvgHash}_{\textit{I}}$ $\text{HistMatch}_{\textit{I}}$ $\text{SNR}_{\textit{A}}$ $\text{SpecDist}_{\textit{A}}$
+Pipeline A 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
+Pipeline B 0.190 0.599 0.095 0.083 0.122 0.276 0.646 0.642 0.249 0.261
+Pipeline C 0.222 0.613 0.137 0.117 0.367 0.347 0.766 0.528 0.247 0.260
+Table 1: Combined evaluation results.
+Plan coherence metrics (left) and modality generation metrics (right) are reported for each pipeline.
+Subscripts denote evaluation domains: T = text, Pl = planning, TS = time series, I = image, and A = audio.
+Scores are averaged over three days; higher scores between Pipeline B and Pipeline C are highlighted.
 Setup and Methodology
@@
 This isolates the specialist model’s execution quality from the orchestrator’s prompt-generation quality.
-Table 1: Plan Coherence Results.
-Scores are averaged over three days, comparing each pipeline’s orchestrator output against a single baseline reference.
-Higher scores between Pipeline B and C are highlighted.
-Pipeline ROUGE-L BERTScore-F1 PlanningLCS PlanningJaccard TimeSeriesDTW
-Pipeline A 1.000 1.000 1.000 1.000 1.000
-Pipeline B 0.190 0.599 0.095 0.083 0.122
-Pipeline C 0.222 0.613 0.137 0.117 0.367
-Table 2: Modality Generation Quality Results.
-Scores are averaged over three days, comparing each pipeline’s specialist model outputs to a high-quality reference generated from the same prompt.
-Higher scores between Pipeline B and C are highlighted.
-Pipeline Image SSIM Image AvgHash Image HistMatch Audio SNR Audio SpecDist
-Pipeline A 1.000 1.000 1.000 1.000 1.000
-Pipeline B 0.276 0.646 0.642 0.249 0.261
-Pipeline C 0.347 0.766 0.528 0.247 0.260
 Results and Analysis
-The quantitative results, summarized in Tables 1 and 2 and visualized in Figure 3, demonstrate GAICo’s value in accelerating AI development and improving system safety.
+The quantitative results, summarized in Table 1 and visualized in Figure 3, demonstrate GAICo’s value in accelerating AI development and improving system safety.
 Orchestrator Performance (Plan Coherence)
-As shown in Table 1, Pipeline A’s orchestrator (GPT-5), serving as the high-quality baseline, expectedly achieved perfect scores (1.000) across all plan coherence metrics.
+As shown in the left side of Table 1, Pipeline A’s orchestrator (GPT-5), serving as the high-quality baseline, expectedly achieved perfect scores (1.000) across all plan coherence metrics.
 In contrast, both Pipeline B (Llama 4) and Pipeline C (Gemini 2.5 Pro) showed significantly lower performance, indicating substantial deviations from the baseline plan.
@@
 Specialist Model Performance (Modality Generation Quality)
-The modality generation quality results (Table 2) reveal the execution fidelity of the specialist models against their respective high-quality references.
+The modality generation quality results (right of Table 1) reveal the execution fidelity of the specialist models against their respective high-quality references.
 Pipeline A’s specialist models achieved perfect scores (1.000) as their outputs were used to define the modality-specific references.
@@
 By separating “Plan Coherence” from “Modality Generation Quality”, we can precisely diagnose failures.
-For instance, the results show that Pipeline B’s poor performance stems from both a weak orchestrator (low scores in Table 1) and less-than-optimal specialist models (Table 2).
+For instance, the results show that Pipeline B’s poor performance stems from both a weak orchestrator (low scores in Table 1 left) and less-than-optimal specialist models (Table 1 right).
 This granular insight allows for targeted optimization, improving the orchestrator’s planning logic or replacing a specific specialist model, which is essential for building trustworthy systems and preventing incidents caused by underperforming components.
@@
 Table 2 summarizes the available examples, which are all publicly accessible in our code repository.
-Table 3: Summary of GAICo Example Notebooks. Those referenced in paper are emphasized. All available at:
-https://github.com/ai4society/GenAIResultsComparator/tree/main/examples.
 Notebook Name Description
@@
 17. example-viz.ipynb Demonstrations of creating custom visualizations for evaluation results.
+Table 2: Summary of GAICo Example Notebooks. Those referenced in this paper are emphasized. All available at:
+https://github.com/ai4society/GenAIResultsComparator/tree/main/examples.
 Conclusion
@@
 GAICo addresses a critical gap in GenAI evaluation by providing the first unified, extensible, and deployed framework for reproducible assessment across diverse data modalities.
-Recent studies have found that even with access to the original source code, many ML experiments cannot be reproduced, mainly due to missing details in documentation and inconsistent evaluation procedures (Lopresti and Nagy 2021).
+Recent studies have found that even with access to the original source code, many ML experiments cannot be reproduced, mainly due to missing details in documentation and inconsistent evaluation procedures (lopresti2021reproducibility).
 GAICo replaces this current landscape of ad-hoc, non-comparable scripts with a standardized approach, integrating metrics for text, structured data (planning, time-series), and multimedia.
@@
 A key priority is to incorporate metrics for fairness, bias, toxicity, and operational statistics, providing developers with a more holistic toolkit for building safer and more responsible AI systems.
-GAICo’s extensible design makes this easy (see UML class diagram, Figure 7).
-Future work also includes exploring developing interactive dashboards or deeper integration with MLOps platforms (like MLflow (Zaharia et al.
-2018)) to enable dynamic exploration and tracking of evaluation results.
+GAICo’s extensible design makes this easy.
+Future work also includes exploring developing interactive dashboards or deeper integration with MLOps platforms (like MLflow (Zaharia_Accelerating_the_Machine_2018)) to enable dynamic exploration and tracking of evaluation results.
 Lastly, further work would also focus on adding general-purpose comparison capabilities for complex structured data (e.g., JSON diffing, graph similarity) and extending metrics to handle multi-turn conversational outputs.

### v4

@@
 GAICo empowers AI researchers and developers to efficiently assess system performance, make evaluation reproducible, improve development velocity, and ultimately build more trustworthy AI systems, aligning with the goal of moving faster and safer in AI deployment.
-Since its release on PyPI in Jun 2025, the tool has been downloaded over 15K times, across versions, by Nov 2025, demonstrating growing community interest.
+Since its release on PyPI in Jun 2025, the tool has been downloaded over 16K times, across versions, by Dec 2025, demonstrating growing community interest.
 Deployed Librarypypi.org/project/GAICo
@@
 However, this rapid integration has outpaced the development of standardized, accessible, and domain-appropriate evaluation tools.
-While standard Natural Language Processing (NLP) metrics like BLEU (papineni-etal-2002-bleu) and ROUGE (lin-2004-rouge) are useful, they are often insufficient for structured outputs like AI plans or time-series forecasts.
+While standard Natural Language Processing (NLP) metrics like BLEU (Papineni et al.
+2002) and ROUGE (Lin 2004) are useful, they are often insufficient for structured outputs like AI plans or time-series forecasts.
 In the absence of a unified tool, developers create bespoke evaluation scripts, leading to a fragmented and non-reproducible landscape that hinders the comparison and iterative improvement of AI systems.
@@
 By providing a consistent interface, GAICo directly supports the iterative workflows common in applied AI development.
-According to a PyPI package statistics source (pepy_gaico), the library has been downloaded over 13,000 times in just over 2 months (Jun-Aug 2025), demonstrating its widespread and growing community interest.
+According to a PyPI package statistics source (PyPI Stats 2025), the library has been downloaded over 16,000 times in just over 6 months (Jun-Dec 2025), demonstrating its widespread and growing community interest.
 Common scenarios for usage of GAICo are evaluating an LLM’s output along a range of metrics (Figure 1) and output of multiple LLMs across a single metric (notebooks ‘example-1’ and ‘example-2’ in Table 2).
@@
 General-purpose AI and Natural Language Processing (NLP) evaluation libraries provide foundational metrics for various tasks.
-For instance, the Hugging Face evaluate library (wolf2019huggingface) offers a wide array of metrics for NLP and related tasks, providing a standardized compute method for individual metrics and integrating well within the Hugging Face ecosystem.
-Similarly, Scikit-learn (scikit-learn) provides fundamental metrics for classification, regression, and clustering, adaptable for certain aspects of text evaluation (e.g., cosine similarity on TF-IDF vectors).
-Libraries like NLTK (bird2009natural) and SpaCy (Honnibal_spaCy_Industrial-strength_Natural_2020) offer foundational text processing capabilities and some basic metrics (e.g., BLEU in NLTK).
+For instance, the Hugging Face evaluate library (Wolf et al.
+2019) offers a wide array of metrics for NLP and related tasks, providing a standardized compute method for individual metrics and integrating well within the Hugging Face ecosystem.
+Similarly, Scikit-learn (Pedregosa et al.
+2011) provides fundamental metrics for classification, regression, and clustering, adaptable for certain aspects of text evaluation (e.g., cosine similarity on TF-IDF vectors).
+Libraries like NLTK (Bird et al.
+2009) and SpaCy (Honnibal et al.
+2020) offer foundational text processing capabilities and some basic metrics (e.g., BLEU in NLTK).
 While indispensable, these libraries primarily offer individual text metrics and lack a unified framework for evaluating diverse, multi-modal GenAI outputs or streamlining multi-model comparisons.
@@
 A key differentiator for GAICo is its design choice to decouple evaluation from LLM inference, positioning it as a post-hoc, reproducible comparison framework.
-This contrasts with general-purpose libraries like Hugging Face evaluate by offering a more integrated workflow and specialized metrics, and with end-to-end frameworks, e.g., Ragas (es2024ragas), DeepEval (Ip_deepeval_2025), that are tightly coupled with LLM APIs for “LLM-as-a-judge” evaluations.
+This contrasts with general-purpose libraries like Hugging Face evaluate by offering a more integrated workflow and specialized metrics, and with end-to-end frameworks, e.g., Ragas (Es et al.
+2024), DeepEval (Ip and Vongthongsri 2025), that are tightly coupled with LLM APIs for “LLM-as-a-judge” evaluations.
 While these integrated frameworks are powerful, their coupling can introduce complexities like API costs, rate-limits, and nondeterminism.
@@
 Evaluation in specialized AI domains often relies on custom tools, leading to fragmented practices. GAICo bridges this gap by integrating relevant metrics into a single, consistent framework.
-Automated Planning: Evaluation in automated planning typically focuses on properties like plan length, cost, or state-space coverage, often requiring specialized planners or validators (e.g., PDDL parsers).
+Automated Planning
+Evaluation in automated planning typically focuses on properties like plan length, cost, or state-space coverage, often requiring specialized planners or validators (e.g., PDDL parsers).
 However, when LLMs generate planning sequences, the output might not strictly adhere to formal planning languages, necessitating metrics that compare the content and order of actions rather than strict plan validity.
 GAICo addresses this with PlanningLCS and PlanningJaccard metrics, which are designed to robustly compare action sequences, including concurrent actions, directly from LLM outputs.
-These are inspired by foundational work on measuring plan diversity and similarity (srivastava2006finding; srivastava2007domain) and sequence analysis (hirschberg1977algorithms), adapted specifically for GenAI evaluation contexts.
-Time-Series Analysis: Standard time-series evaluation metrics include Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), or R-squared for forecasting accuracy; however, when comparing generated time-series shapes or sequences of values from LLMs, metrics like Dynamic Time Warping (DTW) (berndt1994using) become crucial for handling variations in speed or phase.
+These are inspired by foundational work on measuring plan diversity and similarity (Srivastava et al.
+2006, 2007) and sequence analysis (Hirschberg 1977), adapted specifically for GenAI evaluation contexts.
+Time-Series Analysis
+Standard time-series evaluation metrics include Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), or R-squared for forecasting accuracy; however, when comparing generated time-series shapes or sequences of values from LLMs, metrics like Dynamic Time Warping (DTW) (Berndt and Clifford 1994) become crucial for handling variations in speed or phase.
 GAICo integrates TimeSeriesDTW and TimeSeriesElementDiff to provide robust, normalized comparisons of time-series data, regardless of their textual representation.
-Multimedia (Image/Audio): Evaluation of generated images and audio often involves specialized libraries and metrics.
-For images, common metrics include Structural Similarity Index (SSIM) (wang2004image), Peak Signal-to-Noise Ratio (PSNR) (sara2019image), or perceptual hashes (farid2021overview).
-For audio, metrics like Signal-to-Noise Ratio (SNR) (yuan2019signal), Perceptual Evaluation of Speech Quality (PESQ) (rix2001perceptual), or Short-Time Objective Intelligibility (STOI) (taal2010short) are used.
-While underlying libraries (e.g., scikit-image (van2014scikit), Pillow (clark2015pillow), librosa (mcfee2015librosa), scipy (2020SciPy-NMeth)) provide these functionalities, GAICo unifies their application within its framework.
-The rise of multi-modal generative models, such as Meta’s ImageBind (girdhar2023imagebind), further underscores the need for unified evaluation tools capable of assessing outputs across diverse modalities.
+Multimedia (Image/Audio)
+Evaluation of generated images and audio often involves specialized libraries and metrics.
+For images, common metrics include Structural Similarity Index (SSIM) (Wang et al.
+2004), Peak Signal-to-Noise Ratio (PSNR) (Sara et al.
+2019), or perceptual hashes (Farid 2021).
+For audio, metrics like Signal-to-Noise Ratio (SNR) (Yuan et al.
+2019), Perceptual Evaluation of Speech Quality (PESQ) (Rix et al.
+2001), or Short-Time Objective Intelligibility (STOI) (Taal et al.
+2010) are used.
+While underlying libraries (e.g., scikit-image (Van der Walt et al.
+2014), Pillow (Clark 2015), librosa (McFee et al.
+2015), scipy (Virtanen et al.
+2020)) provide these functionalities, GAICo unifies their application within its framework.
+The rise of multi-modal generative models, such as Meta’s ImageBind (Girdhar et al.
+2023), further underscores the need for unified evaluation tools capable of assessing outputs across diverse modalities.
 This allows for consistent, normalized comparison of multimedia outputs alongside text and structured data, bridging the gap between general LLM evaluation and highly specialized domain-specific analysis.
@@
 To balance functionality with accessibility, it uses optional dependencies (e.g., pip install ‘gaico[bertscore]’), minimizing the installation footprint.
-The project’s quality is ensured through a comprehensive test suite (pytest (pytest)), continuous integration (CI) for automated testing and deployment, and pre-commit hooks (precommit_2025) for code consistency.
-Extensive documentation (MkDocs (mkdocs_2025)), including runnable Jupyter notebooks and a full API reference, lowers the barrier to adoption for the applied AI community.
+The project’s quality is ensured through a comprehensive test suite (pytest (Krekel et al.
+2004)), continuous integration (CI) for automated testing and deployment, and pre-commit hooks (Sottile 2014) for code consistency.
+Extensive documentation (MkDocs (Christie 2014)), including runnable Jupyter notebooks and a full API reference, lowers the barrier to adoption for the applied AI community.
 Case Study: Evaluating Composite
@@
 GAICo addresses a critical gap in GenAI evaluation by providing the first unified, extensible, and deployed framework for reproducible assessment across diverse data modalities.
-Recent studies have found that even with access to the original source code, many ML experiments cannot be reproduced, mainly due to missing details in documentation and inconsistent evaluation procedures (lopresti2021reproducibility).
+Recent studies have found that even with access to the original source code, many ML experiments cannot be reproduced, mainly due to missing details in documentation and inconsistent evaluation procedures (Lopresti and Nagy 2021).
 GAICo replaces this current landscape of ad-hoc, non-comparable scripts with a standardized approach, integrating metrics for text, structured data (planning, time-series), and multimedia.
@@
 GAICo’s extensible design makes this easy.
-Future work also includes exploring developing interactive dashboards or deeper integration with MLOps platforms (like MLflow (Zaharia_Accelerating_the_Machine_2018)) to enable dynamic exploration and tracking of evaluation results.
+Future work also includes exploring developing interactive dashboards or deeper integration with MLOps platforms (like MLflow (Zaharia et al.
+2018)) to enable dynamic exploration and tracking of evaluation results.
 Lastly, further work would also focus on adding general-purpose comparison capabilities for complex structured data (e.g., JSON diffing, graph similarity) and extending metrics to handle multi-turn conversational outputs.